# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

### v2

@@ 2 Impossibility Results and Lower Bound
(Impossibility without global communication) It is impossible to solve $k$-Dispersion on TVG for any $n\geq 6$ with the agents having $1$-hop visibility but no global communication even if agents are initially co-located. This impossibility result also holds if agents have infinite memory and knowledge of $n$.
Proof
We prove this impossibility on TVG ring $\mathcal{R}$. Let $k=pn$ many agents be co-located at some node $v$ of ring $\mathcal{R}$, where $p$ is any positive integer. If the agents have 1-hop visibility, then as per our visibility model, the agent can sense the missing edge corresponding to some port $p$. There is no global communication. Therefore, the movement of the agents depends on 1-hop visibility and local communication. Since the algorithm is deterministic, the adversary can precompute the agents computation and accordingly remove an edge from $G$. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movements of agents do not lead to dispersion, the adversary does not need to do anything. If the movements of agents lead to dispersion, we discuss how the adversary intervenes. Assume that at the end of round $r$, the agents successfully achieve dispersion. It means that the dispersion is not achieved at the beginning of round $r$, and after the CCM cycle at round $r$, the agents are in the dispersed configuration. Since $k=pn$, each node of $\mathcal{R}$ should have $p$ many agents in the dispersed configuration. At the beginning of round $r$, there is a node $u$ which has at most $p-1$ agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in ring $\mathcal{R}$. Let $z$ ($\leq p-1$) agents be at node $u$. If the adversary does not delete any edge at the beginning of round $r$, then the following is the computation of the agents according to $\mathcal{A}$: let $x_{1}$ agents move from node $u$ to $u_{1}$, $y_{1}$ agents move from node $u$ to $u_{2}$, $x_{2}$ agents move from node $u_{1}$ to $u$, and $y_{2}$ agents move from node $u_{2}$ to $u$. And accordingly, at the end of round $r$, there are $z+(x_{2}-x_{1})+(y_{2}-y_{1})=p$ agents at node $u$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$ and accordingly deletes an edge. A high-level idea behind the impossibility proof is given in Figure 1. We divide the cases as follows.
  • Case $\bm{x_{2}=0}$ or $\bm{y_{2}=0}$: W.l.o.g. let $x_{2}=0$. In this case, the adversary deletes the edge $(u,\,u_{2})$ at the beginning of round $r$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change. Also, since global communication is not there, agents at node $u_{1}$ can not get the information about the missing edge $(u,\,u_{2})$ and hence the movement of agents at node $u_{1}$ will remain the same in round $r$. But the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Therefore, at the end of round $r$, there are $z-x_{1}^{\prime}<p$ agents at the end of round $r$, where $x_{1}^{\prime}$ is the number of agents moving from node $u$ to $u_{1}$ due to the computation at $u$ based on the removed edge. Therefore, at the end of round $r$, the dispersion configuration is not achieved.
  • Case $\bm{x_{1}=0}$ or $\bm{y_{1}=0}$: W.l.o.g. let $x_{1}=0$. At the end of round $r$, there are $z+x_{2}+(y_{2}-y_{1})=p$ agents at node $u$. It is important to note that $x_{2}>0$ and $y_{2}>0$. Based on the values of $y_{1},\,y_{2}$, we have the following cases.
    $\rightarrow$ Sub-case $\bm{y_{1}=y_{2}:}$ The adversary deletes the edge $(u,\,u_{1})$. Therefore, the 1-hop view of agents at node $u_{2}$ does not change, but the 1-hop view of agents at node $u$ and $u_{1}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{2}$ will remain the same in round $r$. In this case, from node $u$, more than $y_{1}$ agents or less than $y_{1}$ agents can not move to node $u_{2}$; otherwise, node $u_{2}$ contains either less than $p$ agents or greater than $p$ agents. If exactly $y_{1}$ agents move to node $u_{2}$, then the number of agents at node $u$ is $z$, and $z$ is less than $p$. Therefore, the dispersion configuration is not achieved at the end of round $r$.
    $\rightarrow$ Sub-case $\bm{y_{1}\neq y_{2}:}$ The adversary deletes the edge $(u,\,u_{2})$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change, but the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{1}$ will remain the same in round $r$. In this case, if agents move from node $u$ to $u_{1}$, then there are at least $p+1$ agents at node $u_{1}$ due to the view of node $u_{1}$. Therefore, the number of agents at node $u$ is $z+x_{2}$. If $z+x_{2}=p$, then $z+x_{2}=z+x_{2}+(y_{2}-y_{1})\implies y_{1}=y_{2}$. This gives a contradiction. Therefore, node $u$ does not contain $p$ agents at the end of round $r$. Therefore, the dispersion is not achieved in this case as well at the end of round $r$.
  • Case $\bm{x_{1},\,x_{2},\,y_{1},\,y_{2}>0}$: It is important to note that $(x_{2}-x_{1})\leq 0$ and $(y_{2}-y_{1})\leq 0$ are not possible. Therefore, either $(x_{2}-x_{1})>0$ or $(y_{2}-y_{1})>0$. W.l.o.g., let $(y_{2}-y_{1})>0$. In this case, the adversary deletes the edge $(u,\,u_{2})$ at the beginning of round $r$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change, but the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{1}$ will remain the same in round $r$. In this case, if more than $x_{1}$ or less than $x_{1}$ agents move from $u$ to $u_{1}$, then at node $u_{1}$, the number of agents is not $p$ due to the fact the view of node $u_{1}$. Therefore, the number of agents at node $u$ is $z+(x_{2}-x_{1})$. If $z+(x_{2}-x_{1})=p$, then $z+(x_{2}-x_{1})=z+(x_{2}-x_{1})+(y_{2}-y_{1})\implies(y_{2}-y_{1})=0$. This gives a contradiction. Therefore, at the end of round $r$, the number of agents at node $u$ is not $p$. Hence, the dispersion is not achieved at the end of round $r$.
This completes the proof. ∎
[…]

### v3

@@ 3 Related work
[33] It is impossible to solve dispersion of $k\geq 5$ agents on a dynamic graph deterministically with the agents having 1-neighborhood knowledge and unlimited memory, but without global communication.
Theorem 3.2.
[33] It is impossible to solve dispersion of $k\geq 3$ agents on a dynamic graph deterministically with the agents having global communication and unlimited memory, but without 1-neighborhood knowledge.
They provide an asymptotically optimal $\Theta(k)$ round algorithm in that setting. This communication model is strictly stronger than the one we consider. Translating their results into our model, one might infer that dispersion for $k\leq n$ agents in 1-interval connected dynamic graphs requires at least 1-hop visibility and 1-hop communication. This indicates a gap between their sufficient conditions and the necessary conditions in our model. Saxena et al. [24] study on the same model [33] and extend the analysis to $T$-path connected graphs ivivivA dynamic graph $\mathcal{G}$ is $T$-Path Connected for $T\geq 1$, if for all $r\in\mathbb{N}\cup\{0\}$ and for any $u$, $v\in V$, there exists at least one round $i\in[r,r+T-1]$ such that there exists a path between $u$ and $v$ in $\mathcal{G}_{i}$.. Since $T$-path connectivity is weaker than 1-interval connectivity, Theorem 3.1 and Theorem 3.2 are valid in $T$-path connectivity as well. They design an algorithm using 1-hop visibility and global communication, and also show that dispersion becomes impossible in connectivity-time dynamic graphs even when agents have full visibility, global communication, and knowledge of all system parameters. Refer to Table 1 and Table 2 for the results from [33, 24] in our model.
The dynamic graphs in both [33, 24] are weaker than time-varying graphs (TVGs) because they lack an underlying footprint. Yet, their communication assumptions are quite restrictive. In contrast, we generalize the communication model by introducing $l$-hop communication, which captures a spectrum between f-2-f ($l=0$) and global communication ($l=\widetilde{D}$). Their results identify conditions that are sufficient for dispersion with $k\leq n$ agents in weak dynamic models, but they do not establish what is necessary for stronger models like TVGs. They kept this as an open question in their conclusion. Our work fills this gap: we establish necessary conditions for solving $k$-balanced dispersion in temporally connected and $l$-bounded 1-interval connected TVGs. Since TVGs are stronger than the models used in [33, 24], our necessary conditions apply to those as well. While the question of necessary conditions for dispersion in static graphs remains open, we provide a necessary condition to solve $k$-balanced dispersion in TVGs.
Table 1: Necessary conditions to solve $k$-balanced dispersion.
Connectivity Number of agents Initial configuration Necessary conditions
1-interval connected, No footprint, [33] $5\leq k\leq n$ Any undispersed configuration 1-hop visibility, 1-hop communication
Connectivity time, No footprint, [24] $3\leq k\leq n$ Any undispersed configuration Impossible
$T$-path connected, No footprint, [24] $3\leq k\leq n$ Any undispersed configuration 1-hop visibility, 1-hop communication
Temporal connectivity (TVGs) $k=pn+q$; $p,q\in\mathbb{N}\cup\{0\}$, $3\leq q\leq n-1,n\geq 4$ At least one node with $\geq p+2$ agents Impossible
Temporal connectivity $k=pn+q$; $p,q\in\mathbb{N}\cup\{0\}$, $3\leq q\leq n-3,n\geq 6$ Any undispersed configuration Impossible
11-bounded 1-interval connected $k=pn$, $p\in\mathbb{N}$ Any undispersed configuration 1-hop visibility, 1-hop communication
$\ell$-bounded 1-interval connected ($\ell\geq 25$) $k=pn+q$; $p,q\in\mathbb{N}\cup\{0\}$, $5\leq q\leq n-1$, $n\in[5,\lfloor\sqrt{\ell}\rfloor]$ At least one node with $\geq p+2$ agents 1-hop visibility, 1-hop communication
Table 2: Algorithmic results of $k$-balanced dispersion.
Connectivity Number of agents Initial configuration Sufficient condition Time complexity Agent memory
1-interval connected, No footprint [33] $k\leq n$ Any 1-hop visibility, global communication $\Theta(k)$ $\Theta(\log n)$
$T$-path connected, No footprint [24] $k\leq n$ Any 1-hop visibility, global communication, knowledge of $T$ $\Theta(kT)$ $\Theta(\log\max\{k,T\})$
$\ell$-bounded 1-interval connected $k\geq 1$ Any 1-hop visibility, global communication $O(k)$ $O(\log n)$
11-bounded 1-interval connected $k=pn+q$, $p\in\mathbb{N}\cup\{0\}$, $q\in\{1,2\}$ Co-located 0-hop visibility, f-2-f communication, knowledge of $n$ $O(n^{4}\max(\log n,\log p))$ $O(\max\{\log n,\log p\})$
4 Our contribution
In this work, we introduce $k$-balanced dispersion, which is close to the load balancing problem. The details of our contributions are summarized as follows.
  1. 1.
    Let $k=pn+q$; $p,q\in\mathbb{N}\cup\{0\}$, $3\leq q\leq n-1,n\geq 4$. We have shown that if initially there is at least one node in $G$ with $p+2$ agents, the problem of $k$-balanced dispersion is impossible to solve in temporally connected dynamic graphs. This holds even if nodes have infinite storage, and agents have infinite memory, full visibility, global communication, and complete knowledge of both $k$ and $n$. (Theorem 5.1)
  2. 2.
    Let $k=pn+q$; $p,q\in\mathbb{N}\cup\{0\}$, $3\leq q\leq n-3,n\geq 6$. We have shown that if the initial configuration is undispersed, the problem of $k$-balanced dispersion is impossible to solve in temporally connected dynamic graphs. This holds even if nodes have infinite storage, and agents have infinite memory, full visibility, global communication, and complete knowledge of both $k$ and $n$. (Theorem 5.2)
  3. 3.
    Let $k=pn$, $p\in\mathbb{N}$. If the initial configuration is undispersed, then to solve $k$-balanced dispersion in 1-bounded 1-interval connected graphs, agents need 1-hop visibility and 1-hop communication. (Theorem 6.1 and Theorem 6.2)
  4. 4.
    Let $\ell\geq 25$ and $k=pn+q$; $p,q\in\mathbb{N}\cup\{0\}$, $5\leq q\leq n-1$, $n\in[5,\lfloor\sqrt{\ell}\rfloor]$. If initially there is at least one node in $G$ with $p+2$ agents, then to solve $k$-balanced dispersion in $\ell$-bounded 1-interval connected graphs, agents need 1-hop visibility and 1-hop communication. (Theorem 6.3 and Theorem 6.4)
  5. 5.
    We provide an algorithm which solves $k$-balanced dispersion in $\ell$-bounded 1-interval connected graphs with $k\geq 1$ agents in $O(k)$ rounds using $O(\log n)$ bits of memory per agent in the synchronous setting with global communication and 1-hop visibility (Section 7).
  6. 6.
    We provide an algorithm which solves $k$-balanced dispersion in $1$-bounded 1-interval connected graphs with $k\in\{pn+1,pn+2\}$ agents, $p\in\mathbb{N}$, in $O(n^{4}\max(\log n,\log p))$ rounds using $O(\max\{\log n,\log p\})$ bits of memory per agent in the synchronous setting with global communication and 1-hop visibility (Section 8).
Refer to Table 1 for the necessary assumptions to solve $k$-balanced dispersion across different connectivity models, and refer to Table 2 for the sufficient assumptions to solve $k$-balanced dispersion in the same models.
5 Impossibility for temporally connected graphs
In this section, we prove that $k$-balanced dispersion is impossible to solve in temporally connected graphs. We begin by constructing a sequence of static graphs $\mathcal{S}_{\mathcal{G}}=\mathcal{G}_{0},\,\mathcal{G}_{1},\,\ldots$, which is temporally connected graph. Then, we show that dispersion cannot be achieved in any round $r$; that is, it is impossible to solve dispersion in $\mathcal{G}_{r}$ for every $r\in\mathbb{N}\cup\{0\}$. We provide two results: (i) initial configuration has at least one node with more than $\lceil k/n\rceil$ agents, (ii) initial configuration is undispersed and has at most $\lceil k/n\rceil$ agents at every node.
5.1 Impossibility with at least one node with more than $\lceil k/n\rceil$ agents
We consider a clique of size $n\geq 4$ as the footprint $G$ of the dynamic graph $\mathcal{G}$. The number of agents is $k=pn+q$, where $p,q\in\mathbb{N}\cup\{0\}$ and $3\leq q\leq n-1$. At round $r$, $\alpha_{r}(u)$ denotes the number of agents at $u$ at the beginning of round $r$, and $\beta_{r}(u)$ denotes the number of agents at the end of round $r$. At round $r$, we denote nodes as $v_{1}^{r}$, $v_{2}^{r}$, …, $v_{n}^{r}$.
The key idea behind our construction is as follows. At every round $r$, the nodes are divided into two connected components: the first component, say $C_{1}^{r}$, includes at most $2$ nodes, and the second component, say $C_{2}^{r}$, includes the remaining nodes. The construction guarantees that if $C_{1}^{r}$ has a single node, it contains at least $p+2$ agents; and if $C_{1}^{r}$ has two nodes, then it contains at least $2p+3$ agents. In both cases, such construction ensures that there’s a node in $C_{1}^{r}$ with at least $p+2$ agents, which prevents successful dispersion. The construction of the dynamic graph $\mathcal{G}$ is as follows.
Figure 2: The graph $\mathcal{G}_{0}$.
Figure 3: (A) The graph $\mathcal{G}_{r-1}$; (B) the graph $\mathcal{G}_{r}$, where $r$ is an odd round.
Construction of the dynamic graph $\mathcal{G}$: Suppose $k$ agents are placed arbitrarily at nodes at round $r=0$. At round 0, there is one node which has at least one node with at least $\lceil k/n\rceil+1=p+2$ agents. Without loss of generality, let $v_{1}^{0}\in\big{\{}v_{i}^{0}:\alpha_{0}(v_{i}^{0})=\max\{\alpha_{0}(v_{i}^{0}):i\in[1,n]\}\big{\}}$.
  • 1.
    $\bm{r=0}$: Define the set of edges as $X_{1}=\big{\{}\overline{v_{1}^{0}v_{j}^{0}}:j\in[2,n]\big{\}}$. At round $r$, we define $\rho(e,r)$ for every $e\in E$ as follows:
    $\rho(e,r)=\begin{cases}0,&\text{if }e\in X_{1}\\ 1,&\text{if }e\in E\setminus X_{1}\end{cases}$
    The resulting graph has two connected components: the first, say $C_{1}^{0}$, is a clique of size 1, while the second, say $C_{2}^{0}$, is a connected component that is a clique of size n1n-1 (see Fig. 2 for $n=4$).
  • 2.
    $\bm{r=2i-1,i\in\mathbb{N}}$: At the end of round r1r-1, the graph consists of two connected component $C_{1}^{r-1}$ and $C_{2}^{r-1}$, where $C_{1}^{r-1}$ is a clique of size 1, and $C_{2}^{r-1}$ is a clique of n1n-1 nodes (see Fig. 3(A) for $n=4$). Let node $v_{1}^{r-1}$ forms $C_{1}^{r-1}$, and nodes $v_{2}^{r-1}$, $v_{3}^{r-1}$, …, $v_{n}^{r-1}$ forms $C_{2}^{r-1}$. Without loss of generality, let $v_{2}^{r-1}\in\big{\{}v_{i}^{r-1}:\beta_{r-1}(v_{i}^{r-1})=\max\{\beta_{r-1}(v_{i}^{r-1}):i\in[2,n]\}\big{\}}$. Define the set of edges as $X_{2}=\big{\{}\overline{v_{i}^{r-1}v_{j}^{r-1}}:i\in[1,2],\;\text{and }j\in[3,n]\big{\}}$. we define $\rho(e,r)$ for every $e\in E$ as follows:
    $\rho(e,r)=\begin{cases}0,&\text{if }e\in X_{2}\\ 1,&\text{if }e\in E\setminus X_{2}\end{cases}$
    The resulting graph has two connected components: the first, say $C_{1}^{r}$, is a clique of size 2, while the second, say $C_{2}^{r}$, is a connected component that is a clique of size n2n-2 (see Fig. 3(B) for $n=4$).
    Figure 4: (A) The graph $\mathcal{G}_{r-1}$; (B) the graph $\mathcal{G}_{r}$, where $r$ is an even round and $r\geq 1$.
  • 3.
    $\bm{r=2i,i\in\mathbb{N}}$: At the end of round r1r-1, the graph consists of two connected component $C_{1}^{r-1}$ and $C_{2}^{r-1}$, where $C_{1}^{r-1}$ is a clique of two nodes, and $C_{2}$ is a clique of n2n-2 nodes (see Fig. 4(A) for $n=4$). Let nodes $v_{1}^{r-1}$ and $v_{2}^{r-1}$ forms $C_{1}^{r-1}$, and nodes $v_{3}^{r-1}$, …, $v_{n}^{r-1}$ forms $C_{2}^{r-1}$. Without loss of generality, let $v_{1}^{r-1}\in\big{\{}v_{i}^{r-1}:\beta_{r-1}(v_{i}^{r-1})=\max\{\beta_{r-1}(v_{i}^{r-1}):i\in\{1,2\}\}\big{\}}$. Define the set of edges as $X_{3}=\big{\{}\overline{v_{1}^{r-1}v_{j}^{r-1}}:j\in[2,n]\big{\}}$. we define $\rho(e,r)$ for every $e\in E$ as follows:
    $\rho(e,r)=\begin{cases}0,&\text{if }e\in X_{3}\\ 1,&\text{if }e\in E\setminus X_{3}\end{cases}$
    The resulting graph has two connected components: the first, say $C_{1}^{r}$, is a clique of size 1, while the second, say $C_{2}^{r}$, is a connected component that is a clique of size n1n-1 (see Fig. 4(B) for $n=4$).
Lemma 5.1.
The dynamic graph $\mathcal{G}$ in our construction maintains temporal connectivity.
Proof.
Consider any round $r\geq 0$ and any two nodes $u$ and $v$. If $u$ and $v$ are in the same connected component, then $\mathcal{J}(u,v,r)\neq 0$, since the component is a clique and $u$ and $v$ are directly connected. Suppose $u$ and $v$ are not in the same connected component. Without loss of generality, let node $u$ be in clique $C_{1}^{r}$, and node $v$ be in clique $C_{2}^{r}$. We divide the proof into two cases: Case (i) the size of clique $C_{1}^{r}$ is one, and Case (ii) the size of clique $C_{1}^{r}$ is two.
  • 1.
    Case (i): As per the construction at round $r+1$, the adversary selects a node $w$ from $C_{2}^{r}$ and forms a clique $C_{1}^{r+1}$ with nodes $u$ and $w$. If $w=v$, then $\mathcal{J}(u,v,r)\neq 0$, since $C_{1}^{r+1}$ is a clique and $u$ and $v$ are directly connected. If $w\neq v$, then $v$ remains in $C_{2}^{r+1}$. At round $r+2$, the adversary selects a node $w^{\prime}$ from $C_{1}^{r+1}$ and forms a clique $C_{2}^{r+2}$ with $w^{\prime}$ and the nodes of $C_{2}^{r+1}$.
    $\ast$ If $w^{\prime}=u$, $\mathcal{J}(u,v,r)\neq 0$, as $u$ and $v$ are directly connected in the clique $C_{2}^{r+2}$.
    $\ast$ If $w^{\prime}\neq u$, $\mathcal{J}(u,v,r)\neq 0$, since there exists a journey from $u$ to $w^{\prime}$ at round $r+1$ (as nodes $u$ and $w^{\prime}$ are in clique $C_{1}^{r+1}$) and from $w^{\prime}$ to $v$ at round $r+2$ (as nodes $w^{\prime}$ and $v$ are in clique $C_{2}^{r+2}$), i.e., $\mathcal{J}=\{(\overline{uw^{\prime}},r+1),(\overline{w^{\prime}v},r+2)\}.$
  • 2.
    Case (ii): As per the construction at round $r+1$, the adversary selects a node $w$ from $C_{1}^{r}$ and forms a clique $C_{2}^{r+1}$ with $w$ and the nodes of $C_{2}^{r}$. If $w=u$, then $\mathcal{J}(u,v,r)\neq 0$, since $C_{2}^{r+1}$ is a clique and $u$ and $v$ are directly connected. If $w\neq u$, then it becomes Case (i) at round $r+1$. Since Case (i) is proved for every $r\geq 0$, therefore, $\mathcal{J}(u,v,r)\neq 0$.
In both cases, we have shown that $\mathcal{J}(u,v,r)\neq 0$. This completes the proof. ∎
Theorem 5.1.
In the constructed dynamic graph, for every round $r\geq 0$, there exists at least one node that has at least $p+2$ agents. This holds even if nodes have infinite storage, and agents have infinite memory, full visibility, global communication, and complete knowledge of both $k$ and $n$.
Proof.
Suppose there exists a deterministic algorithm $\mathcal{A}$ that solves dispersion in our construction. In our construction, at round $r=0$, node $v_{1}^{0}$ forms component $C_{1}^{0}$ and have at least $p+2$ agents. The remaining nodes $v_{2}^{0},\ldots,v_{n}^{0}$ form a clique (component $C_{2}^{0}$). Since no movement can occur between $C_{1}^{0}$ and $C_{2}^{0}$, the number of agents at node $v_{1}$ remains unchanged at the end of round 0. Therefore, the dispersion is not achieved at round $r=0$.
Without loss of generality, let $v_{2}^{0}$ be the node with the maximum number of agents at the end of round 0 in $C_{2}^{0}$; that is, $\beta_{0}(v_{2}^{0})\geq\beta_{0}(v_{i}^{0})$ for all $i\in[3,n]$. As per our construction at round 1, we create two components: $C_{1}^{1}$ consisting of $v_{1}^{0}$ and $v_{2}^{0}$ (a clique of size 2), and $C_{2}^{1}$ consisting of the remaining nodes $v_{3}^{0},\ldots,v_{n}^{0}$ (a clique of size n2n-2). Let $\beta_{0}(v_{1}^{0})=p+q_{1}$, $q_{1}\geq 2$ since $v_{1}^{0}$ has at least $p+2$ agents. The total number of agents in $C_{2}^{0}$ is $pn+q-(p+q_{1})=p(n-1)+(q-q_{1})$. Let $L=p(n-1)+(q-q_{1})$. Then,
$\beta_{0}(v_{2}^{0})\geq\frac{L}{n-1}$
Without loss of generality, let $\beta_{1}(v_{1}^{0})\geq\beta_{1}(v_{2}^{0})$, we have the following inequality.
$\beta_{1}(v_{1}^{0})\geq\left\lceil\frac{\beta_{0}(v_{1}^{0})+\beta_{0}(v_{2}^{0})}{2}\right\rceil$ (1)
Using $\beta_{0}(v_{1}^{0})=p+q_{1}$ and the bound on $\beta_{0}(v_{2}^{0})$, we derive the following from Eq. 1.
$\beta_{1}(v_{1}^{0})\geq p+\left\lceil\frac{(n-2)q_{1}+q}{2(n-1)}\right\rceil$ (2)
Since $q\geq 3$ and $q_{1}\geq 2$, we have the following inequality
$\beta_{1}(v_{1}^{0})$ $\geq p+\left\lceil\frac{2(n-2)+3}{2(n-1)}\right\rceil=p+\left\lceil 1+\frac{1}{2(n-1)}\right\rceil\implies\beta_{1}(v_{1}^{0})\geq p+2$
Hence, at the end of round $r=1$, there exists a node in $C_{1}^{1}$ with at least $p+2$ agents, so dispersion has not been achieved. At round $r=2$, that node $v_{1}^{0}$ becomes isolated again, and forms a clique $C_{1}^{2}$ of size 1, like $v_{1}^{0}$ in round 0. Thus, the situation repeats. By applying this reasoning inductively, dispersion is never achieved in any round $r\geq 0$ as there is at least one node with $p+2$ agents.
Since this construction is independent of node storage, the agents’ memory, visibility, communication model, or any knowledge, no such algorithm $\mathcal{A}$ can exist. This completes the proof. ∎
5.2 Impossibility with every node with at most $\lceil k/n\rceil$ agents
We consider a clique of size $n\geq 6$ as the footprint $G$ of the dynamic graph $\mathcal{G}$. The number of agents is $k=pn+q$, where $p,q\in\mathbb{N}\cup\{0\}$ and $3\leq q\leq n-3$. At round $r$, $\alpha_{r}(u)$ denotes the number of agents at $u$ at the beginning of the round, and $\beta_{r}(u)$ denotes the number at the end. At round $r$, we denote nodes as $v_{1}^{r}$, $v_{2}^{r}$, …, $v_{n}^{r}$.
The key idea behind our construction remains the same as in Section 5.1, with a minor modification. We ensure that in every round, the dynamic graph consists of two connected components such that either some node has fewer than $p$ agents or some node has at least $p+2$ agents. The construction of the dynamic graph $\mathcal{G}$ is as follows.
Construction of the dynamic graph $\mathcal{G}$: Suppose $k$ agents are placed arbitrarily at nodes at round $r=0$. At round 0, there is at least one node which has at most p1p-1 agents. It is important that if $p=0$, then the undispersed configuration has a node with at least $p+2=2$ agents. Therefore, in this section, we can assume $p\geq 1$. For $p=0$, we can use the result from Section 5.1. So, for the rest of this section, we assume $p\geq 1$. Without loss of generality, let $v_{1}^{0}\in\big{\{}v_{i}^{0}:\alpha_{0}(v_{i}^{0})=\min\{\alpha_{0}(v_{i}^{0}):i\in[1,n]\}\big{\}}$.
  • 1.
    $\bm{r=0}$: Define the set of edges as $X_{1}=\big{\{}\overline{v_{1}^{0}v_{j}^{0}}:j\in[2,n]\big{\}}$. At round $r$, we define $\rho(e,r)$ for every $e\in E$ as follows:
    $\rho(e,r)=\begin{cases}0,&\text{if }e\in X_{1}\\ 1,&\text{if }e\in E\setminus X_{1}\end{cases}$
    The resulting graph has two connected components: the first, say $C_{1}^{0}$, is a clique of size 1, while the second, say $C_{2}^{0}$, is a connected component that is a clique of size n1n-1.
  • 2.
    $\bm{r=2i-1,i\in\mathbb{N}}$: At the end of round r1r-1, the graph consists of two connected component $C_{1}^{r-1}$ and $C_{2}^{r-1}$, where $C_{1}^{r-1}$ is a clique of size 1, and $C_{2}^{r-1}$ is a clique of n1n-1 nodes. Let node $v_{1}^{r-1}$ forms $C_{1}^{r-1}$, and nodes $v_{2}^{r-1}$, $v_{3}^{r-1}$, …, $v_{n}^{r-1}$ forms $C_{2}^{r-1}$. If there is a node with at least $p+2$ agents in $C_{2}^{r-1}$ at the end of round r1r-1, then the adversary considers round r=2i1r=2i-1 as $r=0$, and starts using the construction mentioned in Section 5.1. Since we have already shown in Section 5.1 that at every round $r$, there is a node with at least $p+2$ agents, therefore dispersion is not achieved for every round $r\geq 2i-1$. Otherwise, if there is no node with at least $p+2$ agents in $C_{2}^{r-1}$ at the end of round r1r-1, then it constructs $\mathcal{G}_{r}$ as follows.
    Without loss of generality, let $v_{2}^{r-1}\in\big{\{}v_{i}^{r-1}:\beta_{r-1}(v_{i}^{r-1})=\min\{\beta_{r-1}(v_{i}^{r-1}):i\in[2,n]\}\big{\}}$. Define the set of edges as $X_{2}=\big{\{}\overline{v_{i}^{r-1}v_{j}^{r-1}}:i\in[1,2],\;\text{and }j\in[3,n]\big{\}}$. we define $\rho(e,r)$ for every $e\in E$ as follows:
    $\rho(e,r)=\begin{cases}0,&\text{if }e\in X_{2}\\ 1,&\text{if }e\in E\setminus X_{2}\end{cases}$
    The resulting graph has two connected components: the first, say $C_{1}^{r}$, contains a clique of size 2, while the second, say $C_{2}^{r}$, is a connected component that is a clique of size n2n-2.
  • 3.
    $\bm{r=2i,i\in\mathbb{N}}$: At the end of round r1r-1, the graph consists of two connected component $C_{1}^{r-1}$ and $C_{2}^{r-1}$, where $C_{1}^{r-1}$ is a clique of two nodes, and $C_{2}$ is a clique of n2n-2 nodes. Let nodes $v_{1}^{r-1}$ and $v_{2}^{r-1}$ forms $C_{1}^{r-1}$, and nodes $v_{3}^{r-1}$, …, $v_{n}^{r-1}$ forms $C_{2}^{r-1}$. Without loss of generality, let $v_{1}^{r-1}\in\big{\{}v_{i}^{r-1}:\beta_{r-1}(v_{i}^{r-1})=\min\{\beta_{r-1}(v_{i}^{r-1}):i\in\{1,2\}\}\big{\}}$. Define the set of edges as $X_{3}=\big{\{}\overline{v_{1}^{r-1}v_{j}^{r-1}}:j\in[2,n]\big{\}}$. we define $\rho(e,r)$ for every $e\in E$ as follows:
    $\rho(e,r)=\begin{cases}0,&\text{if }e\in X_{3}\\ 1,&\text{if }e\in E\setminus X_{3}\end{cases}$
    The resulting graph has two connected components: the first, say $C_{1}^{r}$, is a clique of size 1, while the second, say $C_{2}^{r}$, is a connected component that is a clique of size n1n-1.
Lemma 5.2.
The dynamic graph $\mathcal{G}$ in our construction maintains temporal connectivity.
Proof.
The proof is similar to Lemma 5.1. ∎
Theorem 5.2.
In the constructed dynamic graph, $k$-balanced dispersion is not achieved at any round $r\geq 0$. This holds even if nodes have infinite storage, and agents have infinite memory, full visibility, global communication, and complete knowledge of both $k$ and $n$.
Proof.
Suppose there exists a deterministic algorithm $\mathcal{A}$ that solves dispersion in our construction. In our construction, at round $r=0$, node $v_{1}^{0}$ forms component $C_{1}^{0}$ and have at most p1p-1 agents. The remaining nodes $v_{2}^{0},\ldots,v_{n}^{0}$ form a clique (component $C_{2}^{0}$). Since no movement can occur between $C_{1}^{0}$ and $C_{2}^{0}$, the number of agents at node $v_{1}$ remains unchanged at the end of round 0. Therefore, the dispersion is not achieved at round $r=0$.
If at the end of round $r=0$, there exists a node in $C_{2}^{0}$ with at least $p+2$ agents, the adversary treats round $r=1$ as the new starting round $r=0$ and proceeds with the construction described in Section 5.1. As shown in Theorem 5.1, this ensures that in every round $r$, at least one node has at least $p+2$ agents, preventing dispersion for all $r\geq 1$. It remains to handle the case where no node in $C_{2}^{0}$ has at least $p+2$ agents at the end of round 0, i.e., every node has at most $p+1$ agents. We now show that even in this case, dispersion is not achieved at the end of round $r=1$.
Without loss of generality, let $v_{2}^{0}$ be the node with the least number of agents at the end of round 0 in $C_{2}^{0}$; that is, $\beta_{0}(v_{2}^{0})\leq\beta_{0}(v_{i}^{0})$ for all $i\in[3,n]$. As per our construction at round 1, we create two components: $C_{1}^{1}$ consisting of $v_{1}^{0}$ and $v_{2}^{0}$ (a clique of size 2), and $C_{2}^{1}$ consisting of the remaining nodes $v_{3}^{0},\ldots,v_{n}^{0}$ (a clique of size n2n-2). Let $\beta_{0}(v_{1}^{0})=p-q_{1}$, $q_{1}\geq 1$ since $v_{1}^{0}$ has fewer than $p$ agents. The value of $\beta_{0}(v_{2}^{0})$ has two choices, either $\leq p$ or $p+1$, as every node of $C_{2}^{0}$ has at most $p+1$ agents at the end of round $r=0$. We discuss each case for the value of $\beta_{0}(v_{2}^{0})$ one by one as follows.
  • 1.
    If $\beta_{0}(v_{2}^{0})\leq p$, then at the end of round $r=1$ in $C_{1}^{1}$, there is a node which has <p<p agents as $\beta_{0}(v_{2}^{0})\leq p$ and $\beta_{0}(v_{1}^{0})<p$.
  • 2.
    If $\beta_{0}(v_{2}^{0})=p+1$, then at the end of round $r=0$, $\beta_{0}(v_{i}^{0})=p+1$ for every $i\in[3,n]$. Therefore, the total agents in $C_{2}^{0}$ is (p+1)(n1)(p+1)(n-1). The value of $k$ is nothing but the sum of agents in $C_{1}^{0}$ and agents in $C_{2}^{0}$.
    $np+q$ $=(n-1)(p+1)+p-q_{1}\implies q=n-1-q_{1}$
    $\implies n-3$ $\geq n-1-q_{1}\;\;\text{as }\;\;q\leq n-3$
[…]